Enter the value of b that makes these two expressions equivalent:
1.5(5x - 2.3) and 7.5x + b

Answer by ankor@dixie-net.com(22740)

1.5(5x - 2.3) = 7.5x + b
distribute the 1.5
7.5x - 3.45 = 7.5x + b
7.5x - 7.5x - 3.45 = b
b = -3.45